OpenStudy (anonymous):

x^2+12x-27=0 solve by using quadratic formula

OpenStudy (anonymous):

\[x^2+12x-27=0\] \[x^2+12x=27\] \[(x+6)^2=27+36\] \[(x+6)^2=63\] \[x+6=\pm\sqrt{63}\] \[x=-6\pm\sqrt{63}\]

OpenStudy (anonymous):

i realize that was not the quadratic formula, that was 'completing the square' but the formula does the exact same thing only in one long complicated step. you will notice that 12 is even so it is easier to complete the square. the quadratic formula forces a denominator on you (2a) and then you have to simplify. here you do not